are the graphs of the line in the pair parallel? explain y=4x+6
-15x+3y=-45

Found 2 solutions by LinnW, stanbon:
Answer by LinnW(1048)

y=4x+6
-15x+3y=-45
The slope of y=4x+6 is 4. Notice that the form is y = mx + b where m is slope.
To find the slope of -15x+3y=-45 , solve for y
-15x+3y=-45
add 15x to each side
3y = 15x -45
divide each side by 3
y = 5x - 15
The slope of this equation is 5 which does not match the slope of the first
equation, since the slope of the first equation is 4.
Since the slopes are not the same, the lines are not parallel.

Answer by stanbon(75887)

are the graphs of the line in the pair parallel? explain
y=4x+6
slope = 4
==================
-15x+3y=-45
y = 5x - 15
slope = 5
-----
The slopes are different so the lines are not parallel.
-----
Cheers,
Stan H.
==========